# EMOS: Embodiment-aware Heterogeneous Multi-robot Operating System with LLM Agents

### v1

@@ 5 Conclusion
Appendix A Appendix
@@ A.1 Capabilities for Embodiment-aware Reasoning
A.2 Habitat-MAS Benchmark Task Design
As mentioned in the previous section, we carefully designed four challenging tasks to evaluate the embodiment-aware reasoning capabilities of MAS. Detailed description of each task are as follows:
@@ A.2 Habitat-MAS Benchmark Task Design
  • Task 2: Cooperative perception for manipulation. Due to limitations in perception caused by the camera’s position and type, some articulated robots like spot may lack the ability to detect target objects on high shelves, while some arm-less robots like Drone with better camera view may succeed. The heterogeneous robots need to cooperate to acquire a good RGB-D perception of objects for precise manipulation. In this single-floor task, different target objects are placed in positions that are visible for certain robots. We aim to test if the MAS can reason about the robot sensor type and viewpoint and successfully assign appropriate robots to perceive all target objects.
@@
  • Task 3: Collaborative single-floor home rearrangement. As we discussed in the last section, different robots have different manipulation capabilities. Articulated robots are limited to reaching objects within their arm’s workspace. For instance, Stretch is equipped with an arm that can extend farther horizontally, while Fetch has a greater vertical reach, allowing it to grasp higher objects compared to Stretch. This single-floor task involves rearranging objects placed in varying positions, including the ground, high shelves, or a bed center far from a navigable area, requiring the robots with different arm workspaces to understand their availability for different rearrangement targets.
@@
  • Task 4: Multi-Robot, multi-object, multi-floor collaborative rearrangement. This is a comprehensive task that requires complex coordination for collaboration. Within this scenario, several distinct types of robots, aerial, wheeled, and legged robots, must collaborate to perceive and rearrange a large set of objects distributed in various positions across multiple floors. This task combines coordinating different capabilities of heterogeneous robots. Specially, some objects are located on high surfaces, such as cabinets upstairs, requiring advanced perception, manipulation and mobility. Our primary goal is to evaluate the MAS’s ability to optimize task execution by effectively leveraging the unique capabilities of each robot, while balancing token efficiency and time step consumption.
A.3 Extra details about Habitat-MAS benchmark
A.3.1 Sub-goal Definition with PDDL language
@@
Habitat environment Puig et al. (2023) have already integrated a PDDL (McDermott et al., 1998) system for composite task definition and goal evaluation in simulation. The goal of a complex task can be defined by a composite logical expression with primitive predicates and logical operators. Based on the PDDL system, we take the following primitive predicates as the sub-goals in evaluation: 1) Robot_At_Object: This is the first stage in every tasks, robot need to firstly navigate to the nearest navigable points to target objects and then execute the following actions like detect or pick. 2) Robot_At_Receptacle: This is the another type of navigation sub-goal for long-horizon tasks like rearrangement ,for which robots need to navigate to the receptacles before placing objects on the recepcles to complete the final goals of rearrangement. 3) Object_At_Receptacle: This is the final goals of rearrangement tasks. Sometimes robots may be assigned to pick and place the objects beyond their reach, which means it is not enough to just count whether robots can navigate to objects or receptacles. We add this sub-goal to test the ability to reason long-horizon task further explain which robot tend to fail in specific tasks. 4) Robot_Detect_Object: This is for specific for perception tasks, aiming to judge how well the system perform in detcting objects.
A.3.2 Episode generation and verification
@@
The episodic datasets in the benchmark are automatically generated by a cascaded sampling and verification process. We first fliter the eligible tasks scenes from HSSD (Khanna* et al., 2023) and MP3D Chang et al. (2017) datasets that meet the defined requirements. For HSSD scenes, we ensure that the navigable points in the scene are connected and that there are a sufficient number of rooms available for placing receptacles and objects. Additionally, we calculate the navigable mesh(navmesh) and verify that the robot can navigate to the closest navigable point near the object. As for MP3D scenes, we ensure that the scene include a multi-floor setup and that the different floors are connected by stairs which are steep enough to prevent wheeled robot from passing through but feasible for spot robot to navigate.
@@
To sample an episode $E=(L,{\mathbf{P}}^{0},T,G)$, we carefully select receptacles with specific height or width characteristics to accommodate different types of objects as $L$, contributing to the layout diversity of the episodes. The robots’ initial position and joint poses ${\mathbf{P}}^{0}$ are alse carefully initialized to ensure that the robot can complete the assigned tasks including navigation, perception and manipulation based on its capabilities. At the same time, the distance between agent initial positions $R_{i}^{0}$, $R_{j}^{0}$, measured by Euclidean distance $d_{\text{euclidean}}(R_{i}^{0},R_{j}^{0})$, are set to avoid collisions due to being too close to each other. By carefully designing goal state $G$ of each task and generating corresponding task descriptions $T$, we ensure that the task description include all objects in the layout, thereby testing MAS comprehensive reasoning and planning capability for the goal state as thoroughly as possible. In mobility and perception task dataset, the length of the navigation path from the robots’ initial position to target object is ensured not to be excessively long, which would lead to increased navigation time, nor too short, which could lead to one of the robots completing the navigation task too quickly and entering a prolonged waiting state, resulting in unnecessary token consumption. An episode would be trivial if line connecting the robot’s starting point and target object forms a straight line and completely coincides with the navigation path, the robot’s navigation process appears mundane. Therefore, when validating the dataset, we ensure that the ratio of the geodesic distance $d_{\text{geodesic}}(R_{i}^{0},O_{j})$ between robot initial position $R_{i}^{0}$ and target object position $O_{j}$ to the Euclidean distance $d_{\text{euclidean}}(R_{i}^{0},O_{j})$ between the robot’s initial position and the object $P(R_{i}^{0},O_{j})$ defined in Equation 2 is greater than 1, while also ensuring that it is not excessively large to avoid overly complex navigation path.
@@
To verify the episode is meaningful, i.e. to disciminate dummy policies and policy based on embodiment-aware reasoning, we introduce a set of validation criteria to 4 tasks, ensuring that the policy requiring multi-step planning and high-level perception outperforms random dummy policies. In specific:
@@
1)Navigation and multi-floor rearrange The object and receptacle in navigation episode we filtered are not on the same floor. In the meantime, fetch robot and spot robot are initialize on the same floor, which means that if the MAS system assigns tasks correctly, spot robot must be assigned to perform cross-floor navigation or rearrange tasks, while the fetch robot(wheeled) can only be assigned to navigate to target objects on the same floor and iteract with them. Besides objects and receptacles placed on different floor in navigation episode, for objects located on the same floor as the robots’ initial position, we excluded object that could be reached by spot robot, ensuring that they can only be operated by fetch robot, which further reduces the success rate in multi-floor rearrange task for dummy policy.
@@
2) Perception and single-floor rearrange For perception and single-floor manipulation tasks, we assume that we initially have the robot type settings with a pair of robots that are equipped with varying capacities within the same ability. We meticulously designed a comparative selection experiment, where robots with identical settings were tasked with completing opposite tasks using a fixed policy (a predefined task execution sequence, similar to an oracle plan). For example, in the positive task, the Fetch robot rearranges object1 and the Stretch robot rearranges object2, while in the negative task, the roles are reversed: the Stretch rearranges object1 and the Fetch rearranges object2. We filtered episodes where the success rates (0 or 1) in the positive and negative experiments were XORed, thereby identifying episodes that can only be completed under specific arrangements. This approach aims to differentiate a dataset that cannot be completed using random (dummy) policies, to further demonstrate the effectiveness of our approach.
A.4 Special case study
@@
Incorrect code generation caused by hallucinations During large-scale experiments, we identified some specific cases that occasionally lead the system to make incorrect judgements about the environment’s state, particularly when the prompt contains numerical information, especially coordinates, even when we use structured text as the prompt for LLM. The code shown in Figure 7 demonstrate how LLM agents verifies whether the robot can interact with the target object based on the object’s height, horizontal distance between nearest navigable point from object to itself given scene description, robot height and arm workspace (max reach of the robot arm) parsed from URDF. The correct verification method is to calculate whether the Euclidean distance from the robot’s center to target object, when the robot is positioned at a navigable point, falls within the robot’ reach. This is the code logic generated by LLM in most case. However, there are still some cases where the large model fails to correctly understand the spatial relationships, even when we explicitly state through structured text that the calculation of whether the object is within the robot’s reach should be based on horizontal distance and height difference. One common scenario is that the code on the right part only compares the height difference between robot and target object with the radius of the robot arm workspace. In other cases, LLM agent may hallucinate a third coordinate beyond the horizontal distance and height difference to compute distance. Both of these cases can lead to the robot misjudging its manipulation range, ultimately resulting in inappropriate task assignments.
A.5 Extra details about experiments
A.5.1 Token enhancement efficiency
@@
EMOS consumes the most token since we include procedures such as leader assignment, group discussion, self reflection, action execution, etc. These procedures all inevitably query and response with large amount of LLM tokens. To evaluate the extra tokens consumed as we develop our framework resulting in higher success rate, we define the relative success rate, relative token usage, and token efficiency. As shown in Table 2, to judge the token efficiency, EMOS outperforms other ablated methods, When comparing the enhancement in success rate and the rise in the token usage, EMOS achieves more boost in success rate than the token usage compared to all other ablated settings. To judge the token used per success rate, which reflects the tokens used for succeeding in the tasks, EMOS performs the best in this metric. It shows that, as the token usage increase, our EMOS multi-agent system still perform the best in token efficiency.
A.5.2 Discussion on observed phenomena
Besides, there are some interesting phenomena: 1) EMOS w/o. group discussion beats EMOS w/o. robot resume in both navigation task success rate (22.42% $\rightarrow$ 14.14%) and average sub-goal success rate (72.45% $\rightarrow$ 65.27%): It indicates that multi-agent role playing might increase hallucination, especially when robots make decisions with overly complex description. 2) EMOS w/o. numerical performs just as badly as EMOS w/o. robot resume in complex tasks (9.20% $\leftrightarrow$ 12.99% in single-floor rearrangment): It suggests that the LLM agents cannot actually reason about the manipulation limitation with URDF, indicating the importance of calling mathematical functions in our EMOS framework. 3) EMOS w/o. robot resume uses more tokens and steps than EMOS w/o. numerical: It reflects that even though EMOS w/o. robot resume can consume fewer tokens with less textual input, the LLM agents must query involving more steps due to the low success rate in planning.
A.5.3 Evaluation on sub-goal success rate
@@
As a supplement, Table 3 shows the sub-goal success rate of the five settings, providing a detail insight in how well each setting performs in planning and sub-tasks execution. All settings perform relatively well in the Robot_At (the first and the second) sub-goals, which just require the robots to navigate to the target points. By comparing EMOS w/o. robot resume with other settings, as we discussed before, due to not having the summary of mobility description, it can be inferred that robots in this setting tend to fail in multi-floor navigation tasks. It is also notable that compared to EMOS w/o. numerical and EMOS w/o. robot resume, EMOS w/o. group discussion achieves higher sub-goal success rates in navigation sub-goals, while lower in detection sub-goals. This reflect that in this setting, MAS tend to fail in the tasks that reqiure robot to collaborate to complete, which affirm the superiority of the group discussion in our framework, which can indeed reduce the homogenization in multi-agent task planning.

### v2

@@ 5 Conclusion
    Appendix A Appendix: Additional details for EMOS
@@ A.1 Capabilities for Embodiment-aware Reasoning
    A.2 Multi-agent System Design and Communication
    Hierarchical MAS Communication Graph We design the MAS of EMOS following the HMAS-2 framework, which has been proven to be the most efficient LLM-based multi-robot communication framework by Chen et al.  (Chen et al., 2024). In particular, this MAS has one leader LLM agent for high-level task planning and subtask assignment, and several robot LLM agents to provide additional feedback back to the leader LLM agent given the assigned subtasks. The MAS communication graph can be referenced in figure 7, similar to a star topology.
    Figure 7: Communication Graph Design for EMOS This figure show our design on MAS communication graph of EMOS, leader LLM agent decompose the task and assign task to each robot LLM agent while robot LLM agents provide individual feedbacks for leader LLM agent to replan.
    As we discussed in section 3.4 and algorithm 1, hierarchical task planning has 2 stages including centralized group discussion and decentralized action execution. The LLM agent leader will decompose the initial task into subtasks and assign them to different robot agents without embodiment reasoning, which can be referred to by the black arrows in figure 7. Then robot LLM agents will then reason with embodiment awareness through robot resumes themselves, judge whether they themselves can complete the given subtask according to its own embodied capability, and feed it back to the leader agent for replanning referring to the red arrows in figure 7. If each robot agent replies yes, then the task assignment will be executed in the following steps. A more detailed explanation of the two stages of the algorithm is provided.
    Centralized Group Discussion and Task Assignment In the first stage of centralized group discussion, there is a CentralPlanner that generates an initial plan for each robot, and each robot also has an LLM agent that checks its assigned subtask and provides feedback to the central planner in Reflection. With the robot resume composed of general description and numerical details, the robot-dedicated agent could reason its general availability for the assigned task and further check the geometric availability by using mathematical tools. Specifically, by accessing robot resumes in the last section, a robot-dedicated agent is encouraged to perform common-sense reasoning with textual summaries, and generate code with numerical data in both the scene description and the robot resume to perform calculations for spatial-aware reasoning.
    Decentralized Action Execution In the second stage, with the result of the assignment of tasks, each robot-dedicated agent starts to execute its action in parallel. Given a subtask description and action execution history, a robot-dedicated agent controls the current agent by LLM FunctionCall with robot control libraries. These robot control libraries are implemented with ground truth world information, classic robot trajectory planners and inverse kinematics solvers. An agent automatically goes to wait state when it finishes all reasoned actions. When the agent fails to accomplish a subtask but ”believes” to have finished it, it could continue to execute the following planned actions. A side note is that, to study this circumstance of partial failure, we also evalute the task by sub-goals, which will be discussed in the experiments. The agent in the wait state will awake when it receives a new task assignment from a group discussion. For an episode, the task ends when all robots are in the wait state.
    Appendix B Appendix: Additional details for Habitat-MAS
    B.1 Habitat-MAS Benchmark Task Design
    As mentioned in the previous section, we carefully designed four challenging tasks to evaluate the embodiment-aware reasoning capabilities of MAS. Detailed description of each task is as follows:
@@ B.1 Habitat-MAS Benchmark Task Design
    • Task 2: Cooperative perception for manipulation. Due to limitations in perception caused by the camera’s position and type, some articulated robots like spot may lack the ability to detect target objects on high shelves, while some arm-less robots like Drone with better camera view may succeed. The heterogeneous robots need to cooperate to acquire a good RGB-D perception of objects for precise manipulation. In this single-floor task, different target objects are placed in positions that are visible for certain robots. We aim to test whether the MAS can reason about the robot sensor type and viewpoint and successfully assign appropriate robots to perceive all target objects.
@@
    • Task 3: Collaborative single-floor home rearrangement. As we discussed in the last section, different robots have different manipulation capabilities. Articulated robots are limited to reaching objects within their arm’s workspace. For instance, Stretch is equipped with an arm that can extend farther horizontally, while Fetch has a greater vertical reach, allowing it to grasp higher objects compared to Stretch. This single-floor task involves rearranging objects placed in varying positions, including the ground, high shelves, or a bed center far from a navigable area, which requires robots with different arm workspaces to understand their availability for different rearrangement targets.
@@
    • Task 4: Multi-Robot, multi-object, multi-floor collaborative rearrangement. This is a comprehensive task that requires complex coordination for collaboration. Within this scenario, several distinct types of robots, aerial, wheeled, and legged robots, must collaborate to perceive and rearrange a large set of objects distributed in various positions across multiple floors. This task combines the coordination of different capabilities of heterogeneous robots. Specially, some objects are located on high surfaces, such as cabinets upstairs, requiring advanced perception, manipulation and mobility. Our primary goal is to evaluate the MAS’s ability to optimize task execution by effectively leveraging the unique capabilities of each robot, while balancing token efficiency and time step consumption.
    B.2 Sub-goal Definition with PDDL language
@@
    Habitat environment Puig et al. (2023) has already integrated a PDDL system (McDermott et al., 1998) for the definition of composite tasks and the evaluation of the objectives in simulation. The goal of a complex task can be defined by a composite logical expression with primitive predicates and logical operators. Based on the PDDL system, we take the following primitive predicates as the sub-goals in evaluation: 1) Robot_At_Object: This is the first stage in every task, in which the robot needs to firstly navigate to the nearest navigable points to target objects and then execute the following actions like detect or pick. 2) Robot_At_Receptacle: This is another type of navigation sub-goal for long-horizon tasks like rearrangement ,for which robots need to navigate to the receptacles before placing objects on the recepcles to complete the final goals of rearrangement. 3) Object_At_Receptacle: This is the final goal of the rearrangement tasks. Sometimes robots may be assigned to pick and place objects beyond their reach, which means that it is not enough to just count whether robots can navigate to objects or receptacles. We add this sub-goal to test the ability to reason long-horizon tasks further explain which robot tend to fail in specific tasks. 4) Robot_Detect_Object: This is for specific perception tasks, aiming to judge how well the system performs in detecting objects.
    B.3 Episode generation and verification
@@
    The episodic datasets in the benchmark are automatically generated by a cascaded sampling and verification process. We first fliter the eligible scenes from HSSD (Khanna* et al., 2023) and MP3D Chang et al. (2017) datasets that meet the defined requirements. For HSSD scenes, we ensure that the navigable points in the scene are connected and that there are a sufficient number of rooms available for placing receptacles and objects. Additionally, we calculate the navigable mesh(navmesh) and verify that the robot can navigate to the closest navigable point near the object. As for MP3D scenes, we ensure that the scene includes a multi-floor setup and that the different floors are connected by stairs which are steep enough to prevent wheeled robot from passing through but feasible for spot robot to navigate.
@@
    To sample an episode $E=(L,{\mathbf{P}}^{0},T,G)$, we carefully select receptacles with specific height or width characteristics to accommodate different types of objects such as $L$, contributing to the layout diversity of the episodes. The robots’ initial position and joint poses ${\mathbf{P}}^{0}$ are carefully initialized to ensure that the robot can complete the assigned tasks including navigation, perception and manipulation based on its capabilities. At the same time, the distance between the initial positions of the agents $R_{i}^{0}$, $R_{j}^{0}$, measured by Euclidean distance $d_{\text{euclidean}}(R_{i}^{0},R_{j}^{0})$, are set to avoid collisions because they are too close to each other. By carefully designing the goal state $G$ of each task and generating corresponding task descriptions $T$, we ensure that the task description includes all objects in the layout, thereby testing MAS comprehensive reasoning and planning capability for the goal state as thoroughly as possible. In mobility and perception task dataset, the length of the navigation path from the robots’ initial position to target object is ensured not to be excessively long, which would lead to increased navigation time, nor too short, which could lead to one of the robots completing the navigation task too quickly and entering a prolonged waiting state, resulting in unnecessary token consumption. An episode would be trivial if the line connecting the robot’s starting point and target object forms a straight line and completely coincides with the navigation path, the robot’s navigation process appears mundane. Therefore, when validating the dataset, we ensure that the ratio of the geodesic distance $d_{\text{geodesic}}(R_{i}^{0},O_{j})$ between the initial position of the robot $R_{i}^{0}$ and the position of the target object $O_{j}$ to the Euclidean distance $d_{\text{euclidean}}(R_{i}^{0},O_{j})$ between the initial position of the robot and the object $P(R_{i}^{0},O_{j})$ defined in Equation 2 is greater than 1, while also ensuring that it is not excessively large to avoid an overly complex navigation path.
@@
    To verify the episode is meaningful, i.e. to disciminate dummy policies and policy based on embodiment-aware reasoning, we introduce a set of validation criteria to 4 tasks, ensuring that the policy requiring multi-step planning and high-level perception outperforms random dummy policies. In particular:
@@
    1)Navigation and multi-floor rearrange. The object and receptacle in the navigation episode we filtered are not on the same floor. In the meantime, fetch robot and spot robot are initialize on the same floor, which means that if the MAS system assigns tasks correctly, spot robot must be assigned to perform cross-floor navigation or rearrange tasks, while the fetch robot(wheeled) can only be assigned to navigate to target objects on the same floor and iteract with them. Besides objects and receptacles placed on different floor in navigation episode, for objects located on the same floor as the robots’ initial position, we excluded object that could be reached by spot robot, ensuring that they can only be operated by fetch robot, which further reduces the success rate in multi-floor rearrange task for dummy policy.
@@
    2) Perception and single-floor rearrange. For perception and single-floor manipulation tasks, we assume that we initially have the robot type settings with a pair of robots that are equipped with varying capacities within the same ability. We meticulously designed a comparative selection experiment, where robots with identical settings were tasked with completing opposite tasks using a fixed policy (a predefined task execution sequence, similar to an oracle plan). For example, in the positive task, the Fetch robot rearranges object1 and the Stretch robot rearranges object2, while in the negative task, the roles are reversed: the Stretch rearranges object1 and the Fetch rearranges object2. We filtered episodes where the success rates (0 or 1) in the positive and negative experiments were XORed, thereby identifying episodes that can only be completed under specific arrangements. This approach aims to differentiate a dataset that cannot be completed using random (dummy) policies, to further demonstrate the effectiveness of our approach.
    B.4 Detailed Implementation of Robot Low-level Control in Benchmark
    In section 4.1, we generally introduce the robot low-level action implementation in our benchmark as we disable the physics simulation of integrated Pybullet Coumans & Bai (2016–2021). Here we will describe the implementation of each low-level action in detail with non-physics simulation and full observed settings.
    For navigation action navigate_to, as we observed a target object like a apple on the table, the object’s position in world coordinate framework can be calculated. Then we infer the path (series of waypoints) from the start position of the robot to the target position on navmesh through greedy search, and start to force the robot to move towards the next waypoint following the navigation algorithm used in Habitat Puig et al. (2023).
    For arm action including move_arm_to, pick, place, since we have turned off the physical simulation including collision detection, the robotic arm control only requires that after the mobile robot observes the target objects, it first moves to a suitable position near the target position. Subsequently, based on the world coordinates of the object, the joint pose corresponding to the object coordinates is calculated through inverse kinematics, to achieve grasping using the suction grasp.
    It should be noted that low-level control is not our concern, but our framework is easily extensive to introduce other control method like RL policy. Since action strategies are not directly related to embodied task planning, we did not discuss the underlying action strategies in great detail in the article.
    Appendix C Appendix: Additional Experiments
    C.1 Special case study
@@
    Incorrect code generation caused by hallucinations During large-scale experiments, we identified some specific cases that occasionally lead the system to make incorrect judgements about the environment’s state, particularly when the prompt contains numerical information, especially coordinates, even when we use structured text as the prompt for LLM. The code shown in Figure 8 demonstrates how LLM agents verify whether the robot can interact with the target object based on the object’s height, horizontal distance between nearest navigable point from object to itself given scene description, robot height, and arm workspace (max reach of the robot arm) parsed from URDF. The correct verification method is to calculate whether the Euclidean distance from the robot’s center to target object, when the robot is positioned at a navigable point, falls within the robot’s reach. This is the code logic generated by LLM in most cases. However, there are still some cases where the large model fails to correctly understand the spatial relationships, even when we explicitly state through structured text that the calculation of whether the object is within the robot’s reach should be based on horizontal distance and height difference. One common scenario is that the code on the right part only compares the height difference between robot and target object with the radius of the robot arm workspace. In other cases, the LLM agent may hallucinate a third coordinate beyond the horizontal distance and height difference to compute distance. Both of these cases can cause the robot to misjudge its range of manipulation, ultimately resulting in inappropriate task assignments.
    C.2 Token enhancement efficiency
@@
    EMOS consumes the most token since we include procedures such as leader assignment, group discussion, self reflection, action execution, etc. All of these procedures inevitably require a large number of LLM tokens. To evaluate the extra tokens consumed as we develop our framework resulting in a higher success rate, we define the relative success rate, relative token usage, and token efficiency. As shown in Table 2, to judge token efficiency, EMOS outperforms other ablated methods, when comparing the improvement in success rate and the increase in token usage, EMOS achieves a more increased success rate than token usage compared to all other ablated settings. To judge the token used per success rate, which reflects the tokens used to perform the tasks, EMOS performs the best in this metric. It shows that, as the token usage increase, our EMOS multi-agent system still perform the best in token efficiency.
    C.3 Discussion on observed phenomena
    Besides, there are some interesting phenomena: 1) EMOS w/o. group discussion beats EMOS w/o. robot resume in both navigation task success rate (22.42% $\rightarrow$ 14.14%) and average sub-goal success rate (72.45% $\rightarrow$ 65.27%): It indicates that multi-agent role playing might increase hallucination, especially when robots make decisions with overly complex description. 2) EMOS w/o. numerical performs just as badly as EMOS w/o. robot resume in complex tasks (9.20% $\leftrightarrow$ 12.99% in single-floor rearrangment): It suggests that LLM agents cannot actually reason about the manipulation limitation with URDF, indicating the importance of calling mathematical functions in our EMOS framework. 3) EMOS w/o. robot resume uses more tokens and steps than EMOS w/o. numerical: It reflects that even though the EMOS w/o. robot resume can consume fewer tokens with less textual input, the LLM agents must query involving more steps due to the low success rate in planning.
    C.4 Evaluation on sub-goal success rate
@@
    As a supplement, Table 3 shows the sub-goal success rate of the five settings, providing a detailed insight into how well each setting performs in planning and execution of sub-tasks. All settings perform relatively well in the Robot_At (the first and the second) sub-goals, which just require the robots to navigate to the target points. By comparing EMOS w/o. robot resume with other settings, as we discussed before, due to not having the summary of mobility description, it can be inferred that robots in this setting tend to fail in multi-floor navigation tasks. It is also notable that compared to EMOS w/o. numerical and EMOS w/o. robot resume, EMOS w/o. group discussion achieves higher sub-goal success rates in navigation sub-goals, while lower in detection sub-goals. This reflect that in this setting, MAS tend to fail in the tasks that require robot to collaborate to complete, which affirm the superiority of the group discussion in our framework, which can indeed reduce the homogenization in multi-agent task planning.
    C.5 Extra Experiment on Format of Robot Resume
    In order to study which form of robot resume the LLM uses for reasoning with the highest efficiency, we conducted experiments among different formats of robot resumes (natural language, JSON, markdown and XML) for embodied task planning, which are generated using GPT-4o.
    Specifically, we sample 10 episodes from the perception task and evaluate the average success rate of each format. The experimental results can be referred to Table 4
    Table 4: Average Success Rate Using Different Format of Robot Resume
    Format Natural Language JSON Markdown XML
    Avg. Succ Rate 0.3 0.7 0.5 0.6
    Our experiments reveal that structured formats, such as JSON and XML, outperform unstructured formats like natural language in achieving higher success rates for robot resumes. Notably, the success rate increases as the format becomes more structured, which aligns with a key observation in recent research on large language models (LLMs). However, during the experiments, we observed that certain formats, such as Markdown and XML, can induce hallucinations in LLMs. In these cases, the agents do not generate the properly formatted actions, resulting in a 0% success rate. To enable meaningful comparisons, we refined the prompts with minimal modifications to address this issue and produce usable results.
    Based on the observations, the JSON format, used in our frameworks, performs the best. To answer the question about how to generate a better-formatted resume, we suggest using structured formats like JSON format in our frameworks, rather than loosely structured formats like natural language.
    C.6 Extra Experiments on Scalability of EMOS
    Scalability is a crucial aspect of designing a robust framework like EMOS, as it enables seamless integration of new features and ensures that the system can adapt to evolving requirements. While the core functionalities of EMOS have demonstrated effectiveness in Habitat-MAS, we also provide additional experiments to evaluate how well the framework accommodates new capabilities and scales across different tasks and conditions. By testing these attributes, we aim to provide some insight about the scalability of EMOS when extending its scale, which could help the community working on scalable multi-agent systems to understand the system characteristics better.
    Scalability with Robot Number In order to further verify whether EMOS can be applied to systems with increasing number of robots, robot types or task complexity to test the scalability, we conduct experiments by scaling the number of robots performing the same task. In our experiments, we sample 10 episodes from the manipulation task and evaluate performance across different numbers of Fetch robots (2, 4, 6, and 10) to assess communication efficiency and success rate.
    As shown in table 5, we found that as the number scales up, the multi-agent system will face problems like hallucinations (in the setting of 10 agents) and the average success rate will decline. This is as expected since the hallucination problem in LLM is prevalent and it becomes worse with the increase of context length. On one hand, this could be alleviated with more powerful LLM models as we have witnessed in the recent progress of LLM models. In the other hand, designs like hierarchical communication with smaller subgroup discussions and larger group aggregation (similar to delegate meeting) could help solve the scalability problem in multi-agent discussion.
    Table 5: Scaling up with Robot Number
    Num of Robots Succ. Rate Token Usage
    2 80% 48779
    4 60% 73202
    6 70% 93252
    10 50% 151952
    Table 6: Scaling up with Object Number
    Num of Objects Succ. Rate Token Usage
    1 90% 25778
    2 80% 50005
    3 80% 87668
    5 70% 197485
    Scalability with Object Number Considering the scalability problem from another perspective, we conduct a second experiment focusing on increasing task complexity by scaling up the number of objects to manipulate, referring to table 6. While maintaining the fixed number of Fetch robots of (2), we evaluate the system’s performance with varying numbers of objects (1, 2, 3, and 5).
    In the second experiment, which involved scaling up task complexity, we found that our system demonstrates robustness to a certain extent. As the number of objects increases — indicating greater task complexity — the system maintains a relatively high and stable success rate (above 70%). However, the increasing hallucination problem still exists under this setting.